A Complete Bibliography of Publications in British Society for the History of Science Monographs

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 November 2016 Version 1.00

Title word cross-reference

72A [RU84].

accounts [Tha99, Tha10]. America [Out80, Out09]. Annales [Cro94]. Archives [MR88]. art [FJ97]. Australasia [Out80, Out09].

Bacon [RU84]. Beagle [Ste11]. biology [LB95]. Britain [Wyn82]. British [LB95, MR88].

calendar [Out80, Out09]. Chemical [MR88]. chimie [Cro94]. chlorine [CJ07]. Club [Tha99, Tha10]. commentary [RU84]. controversy [CJ07]. correspondence [Pye92]. Cuvier [Out80, Out09].

Dawson [Pye92]. decisions [Wyn82].

E. [LB95]. early [Hun82, Hun94]. Earth [JP79, JP97]. element [CJ07]. environmental [JP79, JP97, Sho93]. essays [JP79, JP97, Sho93]. establishment [Cro94]. Europe [Out80, Out09]. eye [Tha99, Tha10].

Fellows [Hun82, Hun94, Tha99, Tha10]. fight [Tha99, Tha10]. Francis [RU84].

Gallery [FJ97]. Geological [Tha99, Tha10]. Georges [Out80, Out09].

Handlist [MR88]. Hardwick [RU84]. history [FJ97, JP79, JP97, Sho93]. HMS [Ste11]. illustrate [FJ97]. Images [JP79, JP97]. Index [Pye92]. Industry [MR88]. Inquiry [Wyn82]. institution [Hun82, Hun94].

John [Pye92].

Lankester [LB95]. Lavoisier [Cro94]. letters [Out80, Out09]. life [CJ07]. London [Tha99, Tha10]. loss [Ste11].

making [LB95]. manuscript [RU84, Out80, Out09]. materials [Out80, Out09]. McCormick [Ste11]. medicine [CJ07]. meetings [Tha99, Tha10]. modern [LB95]. morphology [Hun82, Hun94].

National [FJ97]. Natural [RU84]. nature [Sho93]. no [Ste11]. nuclear [Wyn82].

philosophy [RU84]. preserved [Out80, Out09]. printed [Out80, Out09].

Rationality [Wyn82]. Ray [LB95]. ritual [Wyn82]. Robert [Ste11]. Royal [Hun82, Hun94].

Science [FJ97, Sho93, CJ07, Cro94]. sciences [JP79, JP97, Sho93]. scientific [Hun82, Hun94, Pye92]. see [Tha99, Tha10]. shadow [Cro94]. Society [Tha99, Tha10, Hun82, Hun94]. source [RU84]. States [Out80, Out09]. summary [Out80, Out09].

technology [CJ07, FJ97]. transcription [RU84]. translation [RU84].

United [Out80, Out09].

voyage [Ste11].

war [CJ07]. William [Pye92]. Windscale [Wyn82]. witness [Tha99, Tha10]. works [FJ97].

References

Chang:2007:ECL

[CJ07] Hasok Chang and Catherine Jackson. An element of controversy: the life of chlorine in science, medicine, technology and war, volume 13 of BSHS monographs. British Society for the History of Science, London, UK, 2007. ISBN 0-906450-01-2 (paperback), 0-906450-15-2. ix + 407 pp. LCCN QD181 .C5 E44 2007.

                                                URL http://books.google.com/
                                                books?id=1tTaAAAAMAAJ;   http:
                                                //catalog.hathitrust.org/api/
                                                volumes/oclc/188588619.html.

Crosland:1994:SLA

[Cro94] Maurice (Maurice Pierre) Crosland. In the shadow of Lavoisier: the Annales de chimie and the establishment of a new science, volume 9 of BSHS monographs / British Society for the History of Science. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1994. ISBN 0-906450-10-1. xii + 354 pp. LCCN QD18.F8 C76 1994. URL

                                                http://books.google.com/books?
                                                id=Z9TaAAAAMAAJ;   http://
                                                catalog.hathitrust.org/api/
                                                volumes/oclc/31887829.html.

Field:1997:SAW

[FJ97] J. V. Field and Frank A. J. L. James. Science in art: works in the National Gallery that illustrate the history of science and technology, volume 11 of BSHS monographs. British Society for the History of Science, Stanford in the Vale, Faringdon, UK, 1997.

ISBN 0-906450-13-6. 110 pp. LCCN ND1460.S34 N38 1997.

Hunter:1982:RSF

[Hun82] Michael Cyril William Hunter. The Royal Society and its Fellows, 1660– 1700: the morphology of an early scientific institution, volume 4 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1982. ISBN 0-906450-03-9. 270 pp. LCCN Q41.L85 H95 1982.

Hunter:1994:RSF

[Hun94] Michael (Michael Cyril William)

Hunter. The Royal Society and its Fellows 1660–1700: the morphology of an early scientific institution, volume 4 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, second edition, 1994. ISBN 0-906450-09-8 (paperback). ix + 291 pp. LCCN Q41.L85 H95 1994.

Jordanova:1979:IEE

[JP79] Ludmilla J. Jordanova and Roy Porter, editors. Images of the Earth: essays in the history of the environmental sciences, volume 1. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1979. ISBN 0-906450-00-4. LCCN QE6 .I48. Proceedings of a conference held at New Hall, Cambridge, April 1977 on new perspectives in the history of geology.

Jordanova:1997:IEE

[JP97] Ludmilla J. Jordanova and Roy Porter, editors. Images of the Earth: essays in the history of the environmental sciences, volume 1 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1997. ISBN 0-906450-12-8. LCCN QE6 .I48 1997. Proceedings of a conference entitled “New Perspectives in the History of Geology” held at New Hall, Cambridge, April 1977.

Lester:1995:RLM

[LB95] Joseph Lester and Peter J. Bowler. E. Ray Lankester and the making of modern British biology, volume 10 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1995. ISBN 0-906450-11-X (paperback). 220 pp. LCCN QH31.L3 L478 1995. Volume incorrectly numbered 9.

Morris:1988:ABC

[MR88] Peter J. T. Morris and Colin A. Russell. Archives of the British Chemical Industry, 1750–1914: A Handlist, volume 6 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1988. ISBN 0-09-064500-6-3. 273 pp. LCCN ???? John Graham Smith, editor.

Outram:1980:LGC

[Out80] Dorinda Outram, editor. The letters of Georges Cuvier: a summary calendar of manuscript and printed materials preserved in Europe, the United States of America, and Australasia, volume 2 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1980. ISBN 0-906450-05-5 (paperback). 3 + 102 pp. LCCN Z6616.C95 O97 Q125.

Outram:2009:LGC

[Out09] Dorinda Outram, editor. The letters of Georges Cuvier: a summary calendar of manuscript and printed materials preserved in Europe, the United States of America and Australasia, volume 2 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 2009. ISBN 0-906450-16-0 (paperback). 102 pp. LCCN ????

Pyenson:1992:ISC

[Pye92] Susan Sheets Pyenson. Index to the scientific correspondence of John William Dawson, volume 7 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1992. ISBN 0-906450-07-1. viii + 275 pp. LCCN Q143.D25 I53 1992.

Bacon:1984:FBN

[RU84] Graham Rees and Christopher Upton, editors. Francis Bacon’s Natural philosophy: a new source: a transcription of manuscript Hardwick 72A, with translation and commentary, volume 5 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1984. ISBN 0-906450-04-7. 197 + 2 pp. LCCN B1198 .R4

1984. URL http://books.google.
com/books?id=fCfXAAAAMAAJ;
http://catalog.hathitrust.org/
api/volumes/oclc/14932617.
html.

Shortland:1993:SNE

[Sho93] Michael Shortland, editor. Science and nature: essays in the history of the environmental sciences, volume 8. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1993. ISBN 0-906450-08-X (paperback). viii + 291 pp. LCCN GF13 .S33 1993. URL http:

//books.google.com/books?id=u_
J-AAAAMAAJ;   http://catalog.
hathitrust.org/api/volumes/
oclc/28748717.html.

Steel:2011:NLR

[Ste11] Emily Steel. He is no loss: Robert McCormick and the voyage of HMS Beagle, volume 14 of BSHS monographs. British Society for the History of Science, London, UK, 2011. ISBN 0-906450-18-7 (paperback). x + 63 pp. LCCN QH31.M33 S74 2011.

Thackray:1999:SFF

[Tha99] John C. (John Christopher) Thackray, editor. To see the Fellows fight: eye witness accounts of meetings of the Geological Society of London and its Club, 1822–1868, volume 12 of BSHS monographs. British Society for the History of Science, Stanford in the Vale, Faringdon, UK, 1999. ISBN 0-906450-14-4 (paperback). xviii + 244 pp. LCCN ????

Thackray:2010:SFF

[Tha10] John C. (John Christopher) Thackray, editor. To see the Fellows fight: eye witness accounts of meetings of the Geological Society of London and its Club, 1822–1868, volume 12 of BSHS monographs. British Society for the History of Science, London, UK, 2010. ISBN 0-906450-17-9 (paperback). xviii + 243 pp. LCCN ????

Wynne:1982:RRW

[Wyn82] Brian Wynne. Rationality and ritual: the Windscale Inquiry and nuclear decisions in Britain, volume 3 of BSHS monographs. British Society for the History of Science, Chalfont St Giles, Bucks, UK, 1982. ISBN 0-906450-02-0 (paperback). x + 222 pp. LCCN HD9698.G72 W96x

1982. URL http://books.google.
com/books?id=eH3bAAAAMAAJ;
http://catalog.hathitrust.org/
api/volumes/oclc/17982399.
html.